A 250cm cube flask of glass filled with sea water weighs 350g. Calculate the density of sea water if empty flask weighs 100g.


1
Expert's answer
2019-09-19T09:48:05-0400

"m_{water} = m - m_{flask}; \\
\rho_{water} = \frac{m_{water}}{V} = \frac{m - m_{flask}}{V} = \frac{350\text{ g}-100\text{ g}}{250\text{ cm}^3} = 1\frac{\text{g}}{\text{cm}^3}."